Circular No. 8078

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS M. L. Sitko, University of Cincinnati; D. K. Lynch, R. W. Russell, and

D. Kim, The Aerospace Corporation; and R. B. Perry, Langley Research Center, NASA, report on 3–13-μm spectroscopy of V838 Mon on Feb. 20.4 UT: “Using the Aerospace Broadband Array Spectrograph System on the NASA Infrared Telescope Facility 3-m telescope, we observed a choppy spectrum in which a quasi-continuum could be fit with a temperature of ∼ 1000 K, with considerable temperature uncertainty. In the 8–13-μm window, a moderately strong silicate emission feature with a well-defined central absorption was apparent. The silicate absorption feature with red and blue emission wings resembled that of ‘NML Cygnus’ (Neugebauer, Martz, and Leighton 1965, Ap.J. 142, 399) reported by Molster et al. (2002, A.Ap. 382, 184) and of H II regions reported by Gillett et al. (1975, Ap.J. 200, 609). Narrowband magnitudes (each ± 0.05) were M [4.5 μm] = 2.5 and N [10.2 μm] = 1.4. There was little discernable change since our Jan. 8.4 measurement with the same instrument and telescope (IAUC 8054).”

SUPERNOVAE 2003ao, 2003ar, 2003as, AND 2003at T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that on the spectra (range 370–750 nm) of several supernovae, obtained by P. Berlind on with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph). A spectrum of SN 2003ao (cf. IAUC 8070), taken on Feb. 22.30 UT, shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 2420 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 5000 km/s. The spectrum is consistent with the plateau phase of a normal type-II supernova. A spectrum of SN 2003ar (cf. IAUC 8072), obtained on Feb. 22.50, shows it to be a type-Ia supernova two to three weeks past maximum. A spectrum of SN 2003as, obtained on Feb. 22.12, shows it to be a type-IIn supernova; the spectrum consists of a very blue continuum with narrow and intermediate velocity-width hydrogen emission lines. A spectrum of SN 2003at (cf. IAUC 8074), obtained on Feb. 22.52, shows it to be a type-II supernova. A weak, narrow Hα emission line from a nearby H II region indicates a recession velocity of 7200 km/s for the host galaxy. Using this value, the expansion velocity derived from the minimum of the Hβ line is 9600 km/s. The spectrum consists of a relatively blue continuum with well-developed P-Cyg lines of hydrogen.

2003 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green